# Private From Whom? Minimal Information Leakage in Auctions

### v1

@@ 3.1 Primitives
What do the bidders and auctioneer learn along the path of play of the mechanism? The auctioneer observes the full sequence of information sets reached and actions taken during the play of the game: For a given protocol $P$ and type profile $\bm{\theta}$, the auctioneer observes the game history $(I^{k}(\bm{\theta}),a^{k}(\bm{\theta}))_{k=1}^{T}$. Let $o_{A}(P,\bm{\theta})$ denote this observed history under protocol $P$ at type profile $\theta$. We say $\bm{\theta}\sim_{(P,A)}\bm{\theta}^{\prime}$ if $o_{A}(P,\bm{\theta})=o_{A}(P,\bm{\theta}^{\prime})$: If the auctioneer cannot distinguish between type profiles $\bm{\theta}$ and $\bm{\theta}^{\prime}$, then they are equivalent under $\sim_{(P,A)}$. To implement $\phi$, the auctioneer must at least be able to distinguish between type profiles that prescribe different outcomes. Letting $\mathcal{P}(\sim)$ denote the partition induced by an equivalence relation $\sim$, this condition can be stated in the language of partitions: $\mathcal{P}(\sim_{(P,A)})$ must be finer than $\mathcal{P}(\sim_{\phi})$ where $a\sim_{\phi}b$ if and only if $\phi(a)=\phi(b)$. [^10] Elements $a,b\in\Theta$ are defined to be in the same cell of $\mathcal{P}(\sim)$ if and only if $a\sim b$. Partition $\mathcal{P}$ is finer than partition $\mathcal{P}^{\prime}$ if for any cell $X$ in $\mathcal{P}$, there exists a cell $Y$ in $\mathcal{P}^{\prime}$ such that $X\subseteq Y$.
@@
What about bidders? Bidder $i$ observes $(\theta_{i},(I_{i}^{k}(\bm{\theta}),a_{i}^{k}(\bm{\theta}))_{k=1}^{T_{i}},\omega_{i})$, which is the history of the information sets at which he has been called to play, the actions he has taken at those information sets, and what she observes about the final outcome, $\omega_{i}\in\Omega_{i}$. [^11] Under this formulation, the bidder can observe their sequence of play, but not calendar time. This is also the definition of an experience, restricted to the end of the game, in li2025obvious. Let $o_{i}(P,\bm{\theta})$ denote bidder $i$’s observation in protocol $P$ when bidders have type profile $\bm{\theta}\in\bm{\Theta}$. We may then define what bidder $i$ learns about other bidders under protocol $P$ by defining the following equivalence relation. We say $\bm{\theta}\sim_{(P,i)}\bm{\theta}^{\prime}$ if $o_{i}(P,\bm{\theta})=o_{i}(P,\bm{\theta}^{\prime})$. That is, the bidder can distinguish between $\bm{\theta}$ and $\bm{\theta}^{\prime}$ in protocol $P$ if and only if they induce different experiences for bidder $i$. Similarly, $\sim_{(P,i)}$ induces a partition $\mathcal{P}(\sim_{(P,i)})$ over $\bm{\Theta}$ for $i$.
3.2 Privacy
Using the above setup, we can now introduce our notions of privacy.
Definition 1 (Bidder Privacy).
@@ 3.2 Privacy
Definition 2 (Auctioneer Privacy).
Protocol $P$ reveals (strictly) more to the auctioneer than protocol $P^{\prime}$ if $\mathcal{P}(\sim_{(P,A)})$ is (strictly) finer than $\mathcal{P}(\sim_{(G^{\prime},A)})$.
Bidder privacy for each $i$ and auctioneer privacy form their own partial orders on the space of all protocols in a given environment (identified with the usual ordering over partitions, restricted to the set of partitions induced by feasible protocols). With the competing objectives of bidder and auctioneer privacy, we can define notions of when a protocol is “better” than another and “best” from a pure privacy perspective.
@@
A protocol $P$ is (strictly) privacy dominated by protocol $P^{\prime}$ if $P$ reveals more information to each bidder than $P^{\prime}$ and $P$ reveals more to the auctioneer than $P^{\prime}$ (with at least one strict).
Definition 4 (Privacy Pareto Frontier).
@@
A protocol $P$ which implements $\phi$ lies on the privacy Pareto frontier for $\phi$ if there does not exist a protocol that implements $\phi$ which strictly privacy dominates $P$.

### v2

@@ 3 Model
What does the auctioneer observe from the protocol? The auctioneer implements the choice rule by repeatedly querying a bidder, observing his action, and moving to the successor information set; this continues until the auctioneer arrives at a terminal history. This embeds an assumption that the auctioneer cannot “commit to forget” the content of a bidder’s message. Hence the auctioneer observes the full sequence of information sets reached and actions taken during the play of the game: For a given protocol $P$ and type profile $\bm{\theta}$, the auctioneer observes the game history $(I^{k}(\bm{\theta}),a^{k}(\bm{\theta}))_{k=1}^{T}$. Let $o_{A}(P,\bm{\theta})$ denote this observed history under protocol $P$ at type profile $\theta$. We say $\bm{\theta}\sim_{(P,A)}\bm{\theta}^{\prime}$ if $o_{A}(P,\bm{\theta})=o_{A}(P,\bm{\theta}^{\prime})$: If the auctioneer cannot distinguish between type profiles $\bm{\theta}$ and $\bm{\theta}^{\prime}$, then they are equivalent under $\sim_{(P,A)}$. To implement $\phi$, the auctioneer must at least be able to distinguish between type profiles that prescribe different outcomes. Letting $\mathcal{P}(\sim)$ denote the partition induced by an equivalence relation $\sim$, this condition can be stated in the language of partitions: $\mathcal{P}(\sim_{(P,A)})$ must be finer than $\mathcal{P}(\sim_{\phi})$ where $a\sim_{\phi}b$ if and only if $\phi(a)=\phi(b)$. [^13] Elements $a,b\in\Theta$ are defined to be in the same cell of $\mathcal{P}(\sim)$ if and only if $a\sim b$. Partition $\mathcal{P}$ is finer than partition $\mathcal{P}^{\prime}$ if for any cell $X$ in $\mathcal{P}$, there exists a cell $Y$ in $\mathcal{P}^{\prime}$ such that $X\subseteq Y$.
@@
What do bidders observe from the protocol? First, the bidder can distinguish some information from the outcome alone. Let $\Omega_{i}$ be a partition of $X$, with cells denoted by $\omega_{i}$, describing the outcomes that bidder $i$ can distinguish from the outcome alone: in the auction setting, we assume that a bidder observes his allocation and payment at the conclusion of the auction. [^14] Suppose $\{(q_{j},t_{j})\}_{j=1}^{N}\in X$ represents an outcome with allocation $q_{i}\in\{0,1\}$ and transfer $t_{i}$ for bidder $i$. Then $\{(q_{j},t_{j})\}_{j=1}^{N}\sim_{\Omega_{i}}\{(q_{j}^{\prime},t_{j}^{\prime})\}_{j=1}^{N}$ if $q_{i}=q_{i}^{\prime}$ and $t_{i}=t_{i}^{\prime}$. This corresponds to bidder $i$ being unable to inherently distinguish between two different outcomes which both result in them losing the auction. Furthermore, each bidder observes his own sequence of communication with the auctioneer. In total, bidder $i$ observes $(\theta_{i},(I_{i}^{k}(\bm{\theta}),a_{i}^{k}(\bm{\theta}))_{k=1}^{T_{i}},\omega_{i})$: the sequence of the information sets at which he has been called to play, the actions he has taken at those information sets, and what she observes about the final outcome, $\omega_{i}\in\Omega_{i}$. [^15] Under this formulation, the bidder can observe their sequence of play, but not calendar time. This is also the definition of an experience, restricted to the end of the game, in Li (2017). Let $o_{i}(P,\bm{\theta})$ denote bidder $i$’s observation in protocol $P$ when bidders have type profile $\bm{\theta}\in\bm{\Theta}$. We may then define what bidder $i$ learns about other bidders under protocol $P$ by defining the following equivalence relation. We say $\bm{\theta}\sim_{(P,i)}\bm{\theta}^{\prime}$ if $o_{i}(P,\bm{\theta})=o_{i}(P,\bm{\theta}^{\prime})$. That is, the bidder can distinguish between $\bm{\theta}$ and $\bm{\theta}^{\prime}$ in protocol $P$ if and only if they induce different experiences for bidder $i$. Similarly, $\sim_{(P,i)}$ induces a partition $\mathcal{P}(\sim_{(P,i)})$ over $\bm{\Theta}$ for $i$.
4 Privacy
We adopt the language of Haupt and Hitzig (2025) to define privacy from the auctioneer. Some privacy violations may be necessary to implement a given $\phi$, but other violations are not: the auctioneer learns information about a bidder that is not pivotal to the outcome.
Definition 1 (Contextual Privacy).
Protocol $P$ produces a contextual privacy violation for bidder $i$ at type profiles $(\theta_{i},\bm{\theta}_{-i}),(\theta_{i}^{\prime},\bm{\theta}_{-i})$ if
Protocols can be partially ordered via inclusion on the type profiles where contextual privacy violations occur. A protocol produces no contextual privacy violations for bidder $i$ if the auctioneer only learns from bidder $i$ exactly what is needed to compute the choice rule. For many choice rules, any protocol produces some contextual privacy violation. However, there are still maximally contextually private protocols; as some violations are inevitable, the set of maximally contextually private protocols is often large, each producing contextual privacy violations at different type profiles.
In contrast, bidder information leakages capture our notion of protecting one bidder’s private information from other bidders. Similar to how the auctioneer must learn some components of bidders’ private information to implement $\phi$, bidders can inherently learn about other bidders’ private information from the outcome of $\phi$. However, information leakages go beyond what bidders can inherently—and inevitably—learn.
Definition 2 (Bidder Information Leakage).
Protocol $P$ produces a bidder information leakage to bidder $i$ at type profiles $(\theta_{i},\bm{\theta}_{-i}),(\theta_{i},\bm{\theta}_{-i}^{\prime})$ if
$\phi(\theta_{i},\bm{\theta}_{-i})\sim_{\Omega_{i}}\phi(\theta_{i},\bm{\theta}_{-i}^{\prime})\text{ but }(\theta_{i},\bm{\theta}_{-i})\nsim_{(P,i)}(\theta_{i},\bm{\theta}_{-i}^{\prime})$
A minimally bidder-informative protocol produces no bidder information leakages to any bidder. For any choice rule $\phi$, one way to attain minimal bidder-informativeness is via a sealed-bid protocol, where each bidder directly reports their type to the auctioneer and the auctioneer chooses the outcome. Such a protocol reveals no information to bidders beyond what they observe of the outcome, because each bidder has only one information set. [^16] Strictly speaking, multiple sealed-bid protocols exist since the auctioneer can query bidders in any arbitrary order. These protocols are all equivalent. In a sealed-bid protocol, the auctioneer perfectly learns bidders’ values, and bidders are minimally informed. As such, sealed-bid protocols are minimally contextually private, foreshadowing our impossibility results in the next section. Sealed-bid protocols also are not the unique minimally bidder-informative protocols: Recall that an ascending second-price auction leaks the same amount of information to bidders as sealed-bid second price auctions. In general, protocols that leak information to bidders are minimally bidder-informative if bidders still learn that information solely from their observations of the outcome.
4.1 Impossibility Theorem
We provide a general condition under which contextual privacy and minimal bidder informativeness conflict: the indistinguishable corners condition.
Definition 3.
A choice rule $\phi$ violates the indistinguishable corners condition for bidder $i$ at type profiles $\{\theta_{i},\theta_{i}^{\prime}\}\times\{\bm{\theta}_{-i},\bm{\theta}_{-i}^{\prime}\}$ if:
  1. 1.
    At $\bm{\theta}_{-i}$, the auctioneer should not distinguish between $\theta_{i}$ and $\theta_{i}^{\prime}$:
  2. 2.
    At $\theta_{i}$, bidder $i$ should not distinguish between $\bm{\theta}_{-i}$ and $\bm{\theta}_{-i}^{\prime}$:
  3. 3.
    At $\bm{\theta}_{-i}^{\prime}$, the auctioneer must distinguish between $\theta_{i}$ and $\theta_{i}^{\prime}$:
Figure 1: Indistinguishable Corners Condition Violation
Figure˜1 illustrates a violation of the indistinguishable corners condition. $\alpha,\beta,\gamma\in X$ represent outcomes of the choice rule $\phi$. The shaded colors represent which outcomes bidder $i$ can distinguish: in particular, $\alpha\sim_{\Omega_{i}}\beta$.
This is a generalization of the corners condition, which requires the stronger condition that $\phi(\theta_{i},\theta_{j},\bm{\theta}_{-ij})=\phi(\theta_{i},\theta_{j}^{\prime},\bm{\theta}_{-ij})$, instead of $\phi(\theta_{i},\theta_{j},\bm{\theta}_{-ij})\sim_{\Omega_{i}}\phi(\theta_{i},\theta_{j}^{\prime},\bm{\theta}_{-ij})$; each outcome is trivially in the same partition as itself. In many cases, there are strictly more violations of the indistinguishable corners condition than the corners condition. In the first-price auction, there are no violations of the corners condition but many violations of the indistinguishable corners condition (Lemma 2). In the second-price auction, there are violations of the corners condition, but also distinct violations of the indistinguishable corners condition (Lemma 4).
Suppose the designer implements some choice rule $\phi$ that violates the indistinguishable corners condition. The driving force behind our impossibility theorem is that to prevent bidders from learning information about other bidders’ types, the auctioneer must ask superfluous questions, learning unnecessary information. For intuition, consider Figure˜1. Suppose bidder $i$ has type $\theta_{i}$. If other bidders have type profile $\bm{\theta}_{-i}^{\prime}$, then the auctioneer must eventually ask bidder $i$ to distinguish type $\theta_{i}$ and $\theta_{i}^{\prime}$, in order for the auctioneer to determine the outcome. This means that if the auctioneer never asks bidder $i$ to distinguish $\theta_{i}$ and $\theta_{i}^{\prime}$, then bidder $i$ understands the type profile is not $(\theta_{i},\bm{\theta}_{-i}^{\prime})$. This is additional information that bidder $i$ obtains from the protocol being run. An auctioneer can instead provide each bidder strictly less information by running a sealed-bid protocol: in a sealed-bid protocol, bidder $i$ cannot distinguish between $\phi(\theta_{i},\bm{\theta}_{-i})$ and $\phi(\theta_{i},\bm{\theta}_{-i}^{\prime})$. This yields a tradeoff: given type profile $\bm{\theta}_{-i}$, either the auctioneer never distinguishes between $\theta_{i}$ and $\theta_{i}^{\prime}$ (which reveals additional information to bidder $i$), or the auctioneer eventually distinguishes between $\theta_{i}$ and $\theta_{i}^{\prime}$ (which produces a contextual privacy violation). Theorem 1 formalizes this insight.
Theorem 1 (Impossibility).
  1. (i)
    $P$ produces no contextual privacy violations for bidder $i$.
  2. (ii)
    $P$ produces no bidder information leakages for any bidder.
The proof of Theorem 1 relies on the following Lemma formalizing the tension between auctioneer and bidder privacy, which follows from the same intuition above. Given some set of auctioneer privacy violations, some instances of bidder information leakages are forced, and vice versa. Lemma 1 is also the workhorse result allowing us to show that certain protocols are on the frontier.
Lemma 1.
Consider $\theta_{i},\theta_{i}^{\prime}\in\Theta_{i}$ and $\bm{\theta}_{-i},\bm{\theta}_{-i}^{\prime}\in\bm{\Theta}_{-i}$. If $(\theta_{i},\bm{\theta}_{-i})\sim_{(P,A)}(\theta_{i}^{\prime},\bm{\theta}_{-i})$ and $(\theta_{i},\bm{\theta}_{-i}^{\prime})\nsim_{(P,A)}(\theta_{i}^{\prime},\bm{\theta}_{-i}^{\prime})$ then $(\theta_{i},\bm{\theta}_{-i})\nsim_{(P,i)}(\theta_{i},\bm{\theta}_{-i}^{\prime})$ and $(\theta_{i}^{\prime},\bm{\theta}_{-i})\nsim_{(P,i)}(\theta_{i}^{\prime},\bm{\theta}_{-i}^{\prime})$. Conversely, if $(\theta_{i},\bm{\theta}_{-i})\sim_{(P,i)}(\theta_{i},\bm{\theta}_{-i}^{\prime})$ then $(\theta_{i},\bm{\theta}_{-i})\sim_{(P,A)}(\theta_{i}^{\prime},\bm{\theta}_{-i})$ if and only if $(\theta_{i},\bm{\theta}_{-i}^{\prime})\sim_{(P,A)}(\theta_{i}^{\prime},\bm{\theta}_{-i}^{\prime})$.
Theorem˜1 implies that a designer need not search for a protocol that preserves both desirable privacy properties; such a protocol does not exist. [^17] The strength of Theorem 1 is that a designer cannot fully preserve any bidder’s privacy from the auctioneer without compromising privacy between bidders. For example, the ascending-join protocol of Haupt and Hitzig (2025) chooses one bidder to protect from the auctioneer, by producing no contextual privacy violations for the chosen bidder. Theorem 1 implies that the ascending-join protocol is not minimally bidder informative. We also use Theorem˜1 to show that several protocols which preserve privacy from the auctioneer may not be preferred by a privacy-conscious designer. In a setting that violates the indistinguishable corners condition, if a protocol $P$ prevents the auctioneer from learning any unnecessary information about some bidder, Theorem˜1 implies we can produce a protocol that is strictly less bidder informative than $P$. This complicates the privacy argument for using such protocols without additionally considering privacy from whom.
4.2 The Privacy Frontier
In light of Theorem˜1, no one protocol is best at protecting privacy from both the auctioneer from bidders. However, protocols can still be partially ordered (via set inclusion) based on the contextual privacy violations and bidder information leakages they result in. If the bidder leakages to bidder $i$ that protocol $P$ produces are a superset of the bidder leakages to bidder $i$ that protocol $P^{\prime}$ produces, then bidder $i$ learns a finer partition of the type space under $P$ than he learns under $P^{\prime}$. A similar definition holds for what the auctioneer learns.
Definition 4.
@@ 4.2 The Privacy Frontier
Definition 5.
Protocol $P$ reveals (strictly) more to the auctioneer than protocol $P^{\prime}$ if $\mathcal{P}(\sim_{(P,A)})$ is (strictly) finer than $\mathcal{P}(\sim_{(P^{\prime},A)})$.
One protocol dominates another from a privacy perspective if it has fewer contextual privacy violations and bidder information leakages. If a protocol is not dominated by any other protocol, it lies on the Privacy Frontier.
@@
A protocol $P$ is (strictly) privacy dominated by protocol $P^{\prime}$ if $P$ reveals more to each bidder than $P^{\prime}$ and $P$ reveals more to the auctioneer than $P^{\prime}$ (with at least one strict).
Definition 7 (Privacy Frontier).
@@
A protocol $P$ which implements $\phi$ lies on the Privacy Frontier for $\phi$ if there does not exist a protocol that implements $\phi$ which strictly privacy dominates $P$.